OpenStudy (mathmath333):

find the remainder of \(\Large\tt \color{black}{30^{72^{87}}}\) divided by 11

ganeshie8 (ganeshie8):

work 72^87 mod 10 first

OpenStudy (mathmath333):

its 8

OpenStudy (mathmath333):

can i apply fermat number here

ganeshie8 (ganeshie8):

yes thats right! may i know how you got 8 ?

OpenStudy (mathmath333):

its 2^87/10

ganeshie8 (ganeshie8):

to be honest, idk the method you're using but it seems very interesting can you teach me quick if possible ? xD

OpenStudy (ikram002p):

its the 2,4,8,6, rule :P

ganeshie8 (ganeshie8):

Oh cycle ?

OpenStudy (mathmath333):

lol its very easy (7*10+2)10 leaves remainder 2^87/10

OpenStudy (mathmath333):

i mean (7*10+2)&87/10 leaves remainder 2^87/10

OpenStudy (mathmath333):

this one (7*10+2)^87/10 leaves remainder 2^87/10

ganeshie8 (ganeshie8):

\[\large \dfrac{72^{87}}{10} = \dfrac{(7*10 + 2)^{87}}{10} = \dfrac{2^{87}}{10}\] ?

ganeshie8 (ganeshie8):

nice :) how do you reduce 2^87 ?

OpenStudy (mathmath333):

its the old remainder theorm i think (ax+b)^n mod a leaves b

ganeshie8 (ganeshie8):

by binomial theorem : (ax+b)^n mod a leaves b^n

OpenStudy (mathmath333):

leaves b^n i mean

ganeshie8 (ganeshie8):

(ax+b)^n = aM + b^n = b^n mod a

ganeshie8 (ganeshie8):

okay whats the trick for reducing 2^87 ?

OpenStudy (mathmath333):

2*2^(86) mod 10= 2^3*16^(21) mod 10 ..and so on u need even power to reduce

OpenStudy (mathmath333):

if u follow the cycle its easiest

ganeshie8 (ganeshie8):

2^87 = 2^3*2^84 = 2^3*(16)^21 = 8*6 = 8 like that ?

OpenStudy (mathmath333):

yes

ganeshie8 (ganeshie8):

Very clever indeed, i use congruences to reduce which is something similar to this but i like your method as it looks more natural xD

OpenStudy (mathmath333):

yes but 1 month earlier u told congruencies are universal

ganeshie8 (ganeshie8):

Ok, so what do we have so far 30^72^87 = 30^(10M+8) right ?

ganeshie8 (ganeshie8):

yes they are still universal, but it requires some theory to be known to appreciate them

OpenStudy (mathmath333):

we use fermat number

ganeshie8 (ganeshie8):

you mean we use Fermat little theorem

OpenStudy (mathmath333):

yes this one

ganeshie8 (ganeshie8):

30^(10M+8) = 30^8*30^(10M) = 30^8*1 mod 11 = ...

OpenStudy (mathmath333):

u mean we can use fermat here ?

ganeshie8 (ganeshie8):

yes since 11 is a prime : \[\large a^{10} \equiv 1 \pmod {11}\]

ganeshie8 (ganeshie8):

see you can't fully avoid congruences :P

OpenStudy (mathmath333):

i always avoid it

OpenStudy (mathmath333):

i use remainder theorm ,but this had a^b^c so i cant use it here

ganeshie8 (ganeshie8):

hmm what do you get after applying fermat little theorem ?

ganeshie8 (ganeshie8):

we got rid of exponent right

ganeshie8 (ganeshie8):

30^72^87 = 30^(10M+8)

OpenStudy (mathmath333):

its 6!!

ganeshie8 (ganeshie8):

show work

OpenStudy (mathmath333):

30^8/11 =8^8/11 =64^4/11 =(-2)^4/11 =16/11 its 5 sry

OpenStudy (mathmath333):

am i right

ganeshie8 (ganeshie8):

Excellent !!

OpenStudy (mathmath333):

so for fermat do i have to always reduce b^c/10 in a^b^c

ganeshie8 (ganeshie8):

thats not fermet, but it depends on what number you're dividing by

ganeshie8 (ganeshie8):

if you're dividing by 13, you reduce b^c/12

ganeshie8 (ganeshie8):

if you're dividing by 19, you reduce b^c/18

ganeshie8 (ganeshie8):

because fermat theorem says this : \[\large a^{p-1}\equiv 1 \pmod p\]

OpenStudy (mathmath333):

ok thanks very much for help

OpenStudy (anonymous):

@ganeshie8 please help me!!